Hankins corporation has 5.4 million shares of common stock outstanding; 290,000 shares of 5.2 percent preferred stock outstanding, par value of $100; and 125,000 5.7 percent semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $72 per share and has a beta of 1.13, the preferred stock currently sells for $103 per share, and the bonds have 20 years to maturity and sell for 103 percent of par. The market risk premium is 6.8 percent, T-bills are yielding 4.3 percent, and the firms tax rate is 23 percent. What is the firms market value capital structure?
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=72*5400000
=388800000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*125000*1.03
=128750000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=103*290000
=29870000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=388800000+128750000+29870000
=547420000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 388800000/547420000
W(E)=0.7102
Weight of debt = MV of Bond/MV of firm
Weight of debt = 128750000/547420000
W(D)=0.2352
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 29870000/547420000
W(PE)=0.0546